OpenStudy (anonymous):

use mathematical induction to show that (7^n)-1 is divisible by 6

OpenStudy (jamesj):

Let P(n) be the statement 7^n - 1 is divisible by 6. Then first of all P(1) is clearly true as 7^1 - 1 = 6. Now suppose that P(k) is true for some k >= 1, we want to show that we can deduce that P(k+1) is true.

OpenStudy (jamesj):

Well, P(k) being true means that 7^k - 1 is divisible by 6 What does that imply for 7^(k+1) - 1 The 'trick' here is finding a way to express this last expression in terms of 7^k - 1. Think about it for a few minutes.

myininaya (myininaya):

so for n=1, it is obiviously true since 6|6 now assume for some k that \[6|(7^k-1)=>7^k-1 =6a, a \in \mathbb{Z} \] so now we to need to show the expression is true for k+1 \[7^{k+1}-1=7^k7-1=7(7^k-\frac{1}{7})=7(7^k-1+\frac{6}{7})=7(7^k-1)+7(\frac{6}{7})\] \[7(6a)+6=6(7a+1)\]

myininaya (myininaya):

but since a is an integer then 7a+1 is an integer so we have that \[6|(7^k-1)\] for all integer k>=1

OpenStudy (anonymous):

so james u mean the thesis will b 7^(k+1)-1

OpenStudy (jamesj):

I'm not exactly sure how you're using the word thesis here. We are trying to show that P(k) ==> P(k+1). I.e., we can deduce the statement P(k+1) from P(k).

hero (hero):

myininaya, where does "a" come from? We started with variable n

myininaya (myininaya):

a is integer that i pulled out of the skies

myininaya (myininaya):

ok i will tell you again if \[6|(7^k-1) => 6a=7^k-1, a \in \mathbb{Z}\]

OpenStudy (jamesj):

We can also do this in an "a"-less way. Note that \[7^{k+1} = 7.7^k - 1 = 7(7^k -1) + 6\] Now we already know something about that expression in brackets, 7^k - 1 ....

myininaya (myininaya):

6|something means there is some integer k such that 6k=something

myininaya (myininaya):

i usually say k but something else was called k so i used a

hero (hero):

James, for less confusion, write: \[7 \dot\ 7^k -1\]

myininaya (myininaya):

i have actually seen 7.7^k written in a paper and it was defined as multiplication

myininaya (myininaya):

well not that exact expression

myininaya (myininaya):

but . was defined as multiplication

myininaya (myininaya):

we can make any symbol to define any operation we want

hero (hero):

Myininaya, stop defending James

myininaya (myininaya):

lol

hero (hero):

lol

myininaya (myininaya):

i like james

OpenStudy (anonymous):

ok james its that all

hero (hero):

I know you do

hero (hero):

You're a biological computer. Biological computers are not allowed to have "favorites" or take sides...: :P

OpenStudy (jamesj):

thanks ... @remainder: do you have your solution?

OpenStudy (anonymous):

thanks every1